OpenStudy (anonymous):

Help with this differential equation please, it won't take my answer in Webwork.

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

did you try the preview answer button? to see if it looks as expected? unless you've used e^ before on ww, you should try exp( )

OpenStudy (anonymous):

Is the answer right though? I'll try typing in preview

OpenStudy (anonymous):

actually it isn't. looked like it was.

OpenStudy (anonymous):

Help me fix it?

OpenStudy (anonymous):

general solution: \[y(t) = c_1 e^{\lambda_1 t} + c_2 e^{\lambda_2 t}\] you found λ 1 and 2 which were 1/5 and 4/5 so we have: \[y(t) = c_1 e^{t/5} + c_2 e^{4t/5}\] the boundary conditions are used to define c and c2. y(0) = 1, when plugged in, leads to c1 + c2 = 1 to use y'(0), we find derivative:\[y'(t) = \frac{ 1 }{ 5 }e^{t/5} + \frac{ 4 }{ 5 }e^{4t/5}\] \[y'(0) = \frac{ 1 }{ 5 }c_1 + \frac{ 4 }{ 5} c_2 = 0 = \frac{ 1 }{ 5 }(c_1 + 4c_2)\] this leads to c1 + 4c2 = 0 combined with c1 + c2 = 1, we can solve for c1 and c2. turns out they are:\[c_1 = -\frac{ 4 }{ 3 }\]\[c_2 = -\frac{ 1 }{ 3 }\] so\[y(t) = \frac{ 4 }{ 3 }e^{t/5} - \frac{ 1 }{ 3 }e^{4t/5}\]

OpenStudy (anonymous):

since y(t) = 4/3e^(t/5) - 1/3e^(4t/5) do we plug in values?

OpenStudy (anonymous):

i forgot to include c1 and c2 in the derivative the first time

OpenStudy (anonymous):

so its y'(t) = 4/3e^(t/5) - 1/3e^(4t/5) or no

OpenStudy (anonymous):

\[y_1 = \frac{ 4 }{ 3 }e^{t/5}\]\[y_2 =- \frac{ 1 }{ 3 }e^{4t/5}\]

OpenStudy (anonymous):

it is, but with c1 and c2, so 4/3 c1 e^(t/5) etc

OpenStudy (anonymous):

when I type those into webwork it says incorrect

OpenStudy (anonymous):

oh wow my bad

OpenStudy (anonymous):

y(t) is the whole thing. y_1 + y_2 the other question has different boundary conditions. just noticed that

OpenStudy (anonymous):

y(t) being y1 on webwork

OpenStudy (anonymous):

I need y1 and y2

OpenStudy (anonymous):

so it should be your screenshot but replace c1 and c2

OpenStudy (anonymous):

I have W(t)

OpenStudy (anonymous):

I never got y1

OpenStudy (anonymous):

y1 isnt: \[\frac{ 4 }{ 3 }e^{t/5} - \frac{ 1 }{ 3 }e^{4t/5}\]?

OpenStudy (anonymous):

it is whats y2 now?

OpenStudy (anonymous):

try solving for c1 and c2 the way i did while i try it too

OpenStudy (anonymous):

using the initial/boundary conditions

OpenStudy (anonymous):

ok but I might get it wrong :(

OpenStudy (anonymous):

i got it. ill give u chance to try though :)

OpenStudy (anonymous):

so whats the first step?

OpenStudy (anonymous):

y(0) = 0 \[y(0) = c_1 e^{0/5} + c_2 e^{4(0)/5} = 0 = c_1 + c_2\] y'(0) = 1\[y'(0) = \frac{ 1 }{ 5 }c_1e^{0/5} + \frac{ 4 }{ 5 }c_2e^{4(0)/5} = 1 = \frac{ 1 }{ 5 }c_1 + \frac{ 4 }{ 5 }c_2\] now that you have 2 equations, you can solve for 1 and 2

OpenStudy (anonymous):

for c1 and c2**

OpenStudy (anonymous):

y2(0) = 0 though

OpenStudy (anonymous):

that's the first line. y(0) = 0

OpenStudy (anonymous):

y'(0) = 1 do you plug the equations back in?

OpenStudy (anonymous):

you plug it in y'(t) which is the derivative of y(t)

OpenStudy (anonymous):

so you said 1/5 c1 + 4/5 c2?

OpenStudy (anonymous):

= 1

OpenStudy (anonymous):

c1 is -4/3 and c2 is -1/3

OpenStudy (anonymous):

1/5e^(t/5)+4/5e^(4t/5) ?

OpenStudy (anonymous):

not in this one. we have\[c_1 + c_2 = 0\]so\[c_1 = -c_2\] then we have\[\frac{ 1 }{ 5 }c_1 + \frac{ 4 }{ 5 }c_2 = 1\] plugging in c1 = -c2:\[-\frac{ 1 }{ 5 }c_2 + \frac{ 4 }{ 5 }c_2 = 1\]leads to\[c_2 = \frac{ 5 }{ 3 }\]plug in c1 = -c2:\[c_1 = -\frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

so whats the final solution?

OpenStudy (anonymous):

\[y_2 = -\frac{ 5 }{ 3 }e^{t/5} +\frac{ 5 }{ 3 }e^{4t/5}\]

OpenStudy (anonymous):

ok I see, some work I just did not understand

OpenStudy (anonymous):

\[y(t) = c_1e^{t/5} + c_2e^{4t/5}\]\[y'(t) = \frac{ 1 }{ 5 }c_1e^{t/5} + \frac{ 4 }{ 5 }c_2e^{4t/5}\] we plugged in the boundary conditions y(0) and y'(0) for the second one y(0) = 0 we plug that in y(t): \[y(0) = 0 = c_1e^{0/5} + c_2e^{4(0)/5}\]since e^0 = 1,\[_1 + c_2 = 0\] same deal for y'(0) = 1

OpenStudy (anonymous):

second to last line: \[c_1 + c_2 = 0\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

glad i could help ^_^